OpenStudy (anonymous):

Find the length of the curve. y=1/6x^3 +1/2x from x=1 to x=2

OpenStudy (anonymous):

\[\int\limits_{1}^{2}F. dS\]

OpenStudy (anonymous):

\[\int\limits_{1}^{2}f(r(t))\left| r'(t) \right| dt\]

OpenStudy (anonymous):

You now need to find the r(t) function

OpenStudy (anonymous):

what is the Rt function?

OpenStudy (anonymous):

its a parameterization. But in your case maybe its easier to use the following formula: \[L = \int\limits_{a}^{2} \sqrt{1 + [f'(x)^{2}]}\]

OpenStudy (anonymous):

where f'(x) its the derivate = (1/2)*(x^2 +1)

OpenStudy (anonymous):

answer = 1.95671

OpenStudy (anonymous):

......